Math 122 (Section 10)  University of South Carolina
D. Meade  Spring 1998

Quiz 1 : Background Information – Math 122 (Spring 1998)

• Personal Information

Name

Phone Number

E-mail Address

Major

Year

• Mathematics and Calculator Background

Yes No Comments

Have you taken the Math Placement Exam? Was your placement code B22 or A41? If not, have you received a “C” or better in Math 111 or 115? Do you have prior experience with a graphing calculator? (If so, which one?) Did you (or will you) purchase a calculator for use this semester?

• Other Information

1. What are your favorite leisure time activities?

2. What are your career goals?

3. Do you have any specific goals or objectives that you hope to fulfill by taking this course?

4. What concerns, if any, do you have about this course?